OpenStudy (anonymous):

Translate to an equation: Y decreased by 3 is identical to three times the number d plus 7. Answer Y - 3 = 3d + 7 3 - Y = 3(d + 7) Y - 3 = 3(d + 7) 3 - Y = 3d + 7

OpenStudy (anonymous):

@iNeedTheHelp

OpenStudy (anonymous):

@yummydum

OpenStudy (anonymous):

y - r = 3D + 7 so A

OpenStudy (anonymous):

@yummydum hows it feel to have your score at 69?

OpenStudy (anonymous):

y - (<--decreased by) 3 = (<-- is equal to) 3(d+7) (<-- 3 times the number d + 7) so its \(y-3=3(d+7)\) so your answer is \(C\)

OpenStudy (anonymous):

@iNeedTheHelp what kind of response are you seeking here?